# Polynomial time sampling from log-smooth distributions in fixed dimension under semi-log-concavity of the forward diffusion with application to strongly dissipative distributions

### v2

@@ 1.2 Our contributions
Denoting $p_{t}^{V}$ the density at time $t$ of the forward OU process started at $\mu\propto e^{-V}$, then under the assumption that $V$ is $C^{2}$ and $L$-smooth, that $\mu$ has finite second moment $m_{2}(\mu)$ and that $-\nabla^{2}\log(p_{t}^{V}),-\nabla^{2}\log(p_{t}^{2V})\succeq-\beta I_{d}$ uniformly over $\mathbb{R}^{d}$, then there exists a stochastic algorithm that outputs a sample $X\sim\hat{p}$ such that $\mathbb{E}[\mathop{\mathrm{KL}}\nolimits(\mu,\hat{p})]\lesssim\epsilon$ in $O(d^{7}L^{d+2}\epsilon^{-2(d+3)}(L+\beta)^{2}(m_{2}(\mu)+d)^{2(d+1)})$ time where $\lesssim$ hides a universal constant as well as log quantities in $d,\epsilon^{-1},L,\beta,m_{2}(\mu)$.
Hence, when the dimension $d$ is fixed, we obtain a polynomial time guarantee. Note that this result pushes one step further the works of Chen et al., 2023b [12], Chen et al., 2023a [10], Conforti et al., [16], Benton et al., [5] who showed that the sampling problem could be transferred into a score approximation problem yet leaving this crucial question open. Using our estimator of the intermediate scores, we show that in fixed dimension, sampling can be framed as a regularity estimate problem: for any class of smooth potentials $V$ such that $\nabla^{2}\log(p_{t}^{V}),\nabla^{2}\log(p_{t}^{2V})$ are bounded from above, we can sample from $\mu\propto e^{-V}$ in polynomial time.
Polynomial time sampling for smooth and strongly dissipative distributions
Deriving classes of potentials $V$ that allow quantitative bounds on $\nabla^{2}\log(p_{t}^{V})$ is an area of research of its own as it allows to recover transport maps from the standard Gaussian to measures of the form $\mu\propto e^{-V}$ with quantitative Lipschitz bounds. Such bounds were originally proven for the class of smooth and strongly convex potentials [27] and later derived for the classes of compactly supported distributions and mixtures of Gaussians with the same variance respectively [35]. In this work, we show that such bounds still hold for smooth potentials that are only assumed to be strongly convex outside some ball. Using Theorem 1, this allows us to recover polynomial complexity bounds for sampling from this class of distributions.
Theorem 2 (Application to smooth and strongly dissipative potentials)
Assume that $V$ is $C^{2}$, $L$-smooth and $\alpha$-strongly convex outside some ball of radius $R$, then there exists a stochastic algorithm that that returns a sample $X\sim\hat{p}$ such that $\mathbb{E}[\mathop{\mathrm{KL}}\nolimits(\mu,\hat{p})]\lesssim\epsilon$ in $O(d^{7}L^{d+8}\epsilon^{-2(d+3)}(R^{2}\kappa^{2}+\frac{d}{\alpha})^{2(d+4)})$ time.
As previously mentioned, this bound exponentially improves upon the $O(e^{32RL^{2}}d\kappa^{2}\epsilon^{-2})$ complexity of ULA [31] and the $O(\frac{e^{40RL^{2}}}{\alpha}\kappa^{3/2}d^{1/2}(d\log(\kappa)+\log(1/\epsilon))^{3/2})$ complexity of MALA [31] in the low dimension and low precision regime. We also show in Section 4 that our results exponentially improve upon Huang et al., [25] and He et al., [22].
1.3 Discussion on concurrent work
In the very recent paper Chak, [8], polynomial time guarantees were also derived for the task of sampling from a distribution that is log-smooth and strongly log-concave outside of a ball that is centered at 00, yet with better exponents than ours. To achieve this, the author relies on IS with a carefully chosen proposal distribution $\nu$: a cutoff parameter $M$ is chosen such that the proposal distribution $\nu$ matches with $\mu\propto e^{-V}$ whenever $V(x)\geq M$ and is flat everywhere else. In Theorem 2.3, such a proposal $\nu$ with matching tails is proven to be a suitable choice for IS. Furthermore, when $M$ is sufficiently large, it is shown in Proposition 5.1 therein that the proposal $\nu$ is log-concave and can thus be safely sampled from using ULA or MALA for instance. However, the cutoff value of $M$ that ensures log-concavity requires estimates on the log-smoothness $L$ and the radius $R$ which may not be available in practice: we discuss in Appendix C a simple, one dimensional example where $R$ depends on the localization of the modes which are unknown a priori. In contrast, our method does not require the knowledge of any of these constants. Furthermore, our method still provides quantitative guarantees under Assumption 1 which is more general than strong log-concavity outside a ball as shown in Section 4.
2 Preliminaries
In this section, we introduce the reverse diffusion framework and review the main existing bounds on how the resulting sampling error is affected by the error on the intermediate scores. Then, we show that these intermediate scores can be re-written as a ratio of expectations and detail how this re-writing differs from previous works.
@@ 2.1 Reverse diffusion: from sampling to score estimation
One of the core tool for sampling from $\mu\propto e^{-V}$ is the Langevin diffusion process $\mathrm{d}X_{t}=–\nabla V(X_{t})\mathrm{d}t+\sqrt{2}\mathrm{d}B_{t}\,,$ where $B_{t}$ is the standard Brownian motion. When discretized with a Euler-Maruyama scheme, this forward equation becomes the well-known Underdamped Langevin Algorithm (ULA, [40]) which is readily implementable whenever we can access $\nabla V$. One of the fundamental questions to determine the quality of the sample output by ULA is the speed of convergence of Langevin towards the equilibrium $\mu$. When $\mu$ is $\alpha$-strongly log-concave, and more generally when $\mu$ verifies an $\alpha$-log-Sobolev inequality [4], exponential convergence in KL occurs at an $\alpha$ rate [46]. Yet, if $\mu$ is multimodal, strong log-concavity no longer holds and the log-sobolev constant $\alpha$ is either no longer defined either scales exponentially with the distance between the modes [11, 9]. This exponential dependence is in correspondence with a phenomenon known as metastability where the process $X_{t}$ gets stuck in a local mode of $\mu$. Reverse diffusion methods emerged as an alternative to Langevin based samplers in order to overcome metastability and were first introduced to the ML community in Song and Ermon, [45]. Instead of targeting $\mu$, one targets the standard Gaussian measure with a Langevin diffusion initialized at $\mu$. We thus obtain the so-called forward process
$\begin{cases}&\mathrm{d}X_{t}=-X_{t}\mathrm{d}t+\mathrm{d}B_{t}\,,\\ &X_{0}\sim\mu\,,\end{cases}$ (1)
which corresponds to the standard Ornstein-Uhlenbeck (OU) process initialized at $\mu$. Note that since we are now targeting the standard Gaussian which is $1$-strongly log-concave, the resulting process converges exponentially fast to the equilibrium. In order to sample from $\mu$ we consider the semi-discretized backward process: given a horizon $T$ that we discretize as $0=t_{0}\leq t_{1}\leq\cdots t_{N-1}\leq t_{N}=T$, the discretized process reads
@@
with $Y_{0}\sim\pi$ and where all iterations can be solved in closed form. Because the forward process Eq. 1 converges exponentially fast, we can expect the initialization error $Y_{0}\sim\pi$ instead of $Y_{0}\sim p_{T}$ to be small for a large time $T$. Furthermore, if the proxies $s_{t_{k}}$ are sufficiently accurate, one can expect that the process output by the approximate scheme Eq. 3 has a distribution that is close to the target $\mu$. Over the past three years, several works provided quantitative bounds of the error induced by the discretization, the use of an approximate score and the initialization error with respect to different divergences and under various assumptions [6, 28, 12, 10, 16]. Yet, we shall rely exclusively on the following theorem as they are the most suited to our framework.
Theorem 3 (Chen et al., 2023a [10])
Assume that $\mu$ is $L$-log-smooth and has finite second order moment $m_{2}(\mu)=\int\|x\|^{2}\mathrm{d}\mu(x)$ and choose the following (backward) discretization: $t_{k-1}=t_{k}-c\min(\max(t_{k},1/L),1)$ for $k\geq 2$ with $c=(\log(L)+T)/N$. Then, if $c$ verifies $c\leq\frac{1}{2d}$, denoting $p$ the distribution of the sample $Y_{T}$ output by Eq. 3, it holds that
$\mathop{\mathrm{KL}}\nolimits(\mu,p)\lesssim(d+m_{2}(\mu))e^{-T}+\sum_{k=1}^{N}(t_{k}-t_{k-1})\|\nabla\log(p_{t_{k}})-s_{t_{k}}\|_{L^{2}(p_{t_{k}})}^{2}+\frac{d^{2}(\log(L)+T)^{2}}{N}\,,$
where $\lesssim$ hides a universal constant. If we make the stronger assumption that $p_{t}^{V}$ remains $L$-log smooth along the trajectory, then it holds using the constant step size $t_{k}=kT/N$ that
$\mathop{\mathrm{KL}}\nolimits(\mu,p)\lesssim(d+m_{2}(\mu))e^{-T}+\sum_{k=1}^{N}(t_{k}-t_{k-1})\|\nabla\log(p_{t_{k}})-s_{t_{k}}\|_{L^{2}(p_{t_{k}})}^{2}+\frac{dT^{2}L^{2}}{N}\,,$
@@
where $\lesssim$ hides a universal constant.
In the case where we have an a priori bound on the log-smoothness of $\mu$, we can use the first statement of the theorem that yields a tighter bound in fixed dimension throught the use of a varying step-size depending on $L$. If no such prior knowledge is available, we can use a constant step size at the cost of a stronger assumption and a weaker bound.
Remark 4
In its original statement, Chen et al., 2023a [10, Theorem 5] assumes that
$\frac{1}{T}\sum_{k=1}^{N}(t_{k}-t_{k-1})\|\nabla\log(p_{t_{k}})-s_{t_{k}}\|_{L^{2}(p_{t_{k}})}^{2}\leq\epsilon^{2}\,,$
and directly provides the upper bound with the $T\epsilon^{2}$ term instead. Yet, an inspection of their Proposition 8 shows that we can safely use $\sum_{k=1}^{N}(t_{k}-t_{k-1})\|\nabla\log(p_{t_{k}})-s_{t_{k}}\|_{L^{2}(p_{t_{k}})}^{2}$. Similarly, it assumes that $c$ should be lower than $\frac{1}{Kd}$ with $K$ some universal constant. Yet a careful inspection of Lemma 13 and Lemma 17 shows that we can take $K=2$.
@@
The previous theorems show that under mild assumptions that notably allow for multimodality, the problem of sampling from $\mu$ can be transferred into a score approximation problem along the forward process. In the next subsection, we build an estimator for these intermediate scores that is tractable given the knowledge of the unnormalized density $\mu\propto e^{-V}$, and show how it differs from previous works.
2.2 Construction of our estimator
The key observation to derive an estimator of the intermediate scores is that the forward process Eq. 1 is nearly available in closed form. Indeed, Eq. 1 integrates to $X_{t}=e^{-t}X_{0}+B_{1-e^{-2t}}\,,$ where $B_{s}$ is the standard $d$-dimensional Brownian motion evaluated at time $s$. In particular, conditionally on $X_{0}$, $X_{t}$ has a normal distribution $X_{t}|X_{0}\sim\mathcal{N}(e^{-t}X_{0},(1-e^{-2t})I_{d})$. By Bayes rule, the distribution $p_{t}^{V}$ of the forward process $X_{t}$ initialized at $\mu\propto e^{-V}$ thus integrates up normalizing constants to $p_{t}^{V}(z)\propto\int e^{-\frac{\|e^{-t}x-z\|^{2}}{2(1-e^{-2t})}}e^{-V(x)}\mathrm{d}x\,.$ We thus recover that the score of the forward process re-writes as $\nabla\log(p_{t}^{V})(z)=\ \mathbb{E}_{Y\sim q_{t,z}}\biggl{[}\frac{z-e^{-t}Y}{1-e^{-2t}}\biggr{]}\,,$ where we denoted $q_{t,z}(x)\propto e^{-\frac{\|e^{-t}x-z\|^{2}}{2(1-e^{-2t})}}e^{-V(x)}$. As noted in Huang et al., [24], Huang et al., [25] and He et al., [22], if samples $y_{i}\sim q_{t,z}$ are available, one can recover an efficient empirical estimator of the score given by $\frac{-1}{n(1-e^{-2t})}\sum_{i=1}^{n}y_{i}$ ready to be plugged in the reverse diffusion process Eq. 3. However, by doing so, the problem of sampling from $e^{-V}$ is simply transferred to the problem of sampling from $q_{t,z}$, which becomes almost as hard as $t$ grows. In Huang et al., [24], the intermediate sampling problems are solved using the ULA algorithm and in He et al., [22], the authors use a rejection sampling scheme which is similar in spirit to importance sampling. Hence, it is not surprising that their final upper-bounds do not improve over standard methods as they employ them to solve intermediate problems that are progressively as hard as the original problem. Instead, we observe that after applying the change of variable $y_{t}=ze^{-t}-x$ both on the numerator and on the denominator, we obtain a score formula which re-writes as a ratio of expectations under standard Gaussians $\nabla\log(p_{t}^{V})(z)=\frac{-1}{1-e^{-2t}}\frac{\mathbb{E}[Y_{t}e^{-V(e^{t}(z-Y_{t}))}]}{\mathbb{E}[e^{-V(e^{t}(z-Y_{t}))}]}\,,$ with $Y_{t}\sim\mathcal{N}(0,(1-e^{-2t})I_{d})$. Unlike the previous identity, which involves an expectation under $q_{t,z}$, the expectations above are computationally cheap as we can easily sample from $Y_{t}$. While conventional statistical wisdom would suggest both the numerator and the denominator to be estimated with independent samples, we voluntarily choose to correlate them and implement instead

### v3

@@ 3.1 Main result
[Main result, informal] Suppose that Assumption 1 and  2 hold. Then, for all $\epsilon>0$, there exists a stochastic algorithm whose parameters only depend on $\epsilon$ (and not on the parameters of the problem), that outputs a sample $X\sim\hat{p}$ such that $\mathbb{E}[\mathop{\mathrm{KL}}\nolimits(\mu,\hat{p})]\lesssim\epsilon\beta^{d+3}(b+d)/a^{2}$ in $O(\mathrm{poly}(\epsilon^{-d}))$ queries to $V$, where $\lesssim$ hides a universal constant as well as log quantities in $d,\epsilon^{-1},\beta,a,b$.
In particular, when $d$ is fixed, this algorithm can output a sample from a distribution that is $\epsilon$-close to $\mu$ in expected $\mathop{\mathrm{KL}}\nolimits$ in $\mathrm{poly}((b+1)/a,\beta,\epsilon^{-1})$ running time.
Algorithm Assumptions Oracle Complexity (Total Variation)
ULA [Ma et al., 2019] $((\alpha+L)\mathbf{1}_{\|x\|\geq R}-L)I_{d}\preceq\nabla^{2}V(x)\preceq LI_{d}$ $\nabla V$, $L$ $O(e^{16LR^{2}}(L/\alpha)^{2}d\epsilon^{-1})$
Proposal-based [He and Zhang, 2025] $\|\nabla^{2}V\|\leq L$, $m_{2}\leq M$ $V$, $\nabla V$, $M$, $L$ $O((LM\epsilon^{-1})^{O(d)})$
RD + ULA [Huang et al., 2024b] $\|\nabla^{2}\log(p_{t})\|\leq L$, $m_{2}<\infty$ $\nabla V$ $e^{O(L^{3}\log^{3}((Ld+m_{2}/\epsilon)))}$
RD + Rejection Sampling [He et al., 2024] $V(x)-V(x^{*})\leq L\|x-x^{*}\|^{2}$, $m_{2}<\infty$ $V$, $V^{*}$ $O(L^{d/4}\epsilon^{-d/2}e^{(L\|x^{*}\|^{2}+m_{2})/2})$
RD + Self-normalized IS (ours) $\nabla^{2}V\preceq\beta I_{d},\;\langle\nabla V(x),x\rangle\geq a\|x\|^{2}-b$ $V$ $O(\sqrt{d}\beta^{(d+3)/2}\sqrt{b+d}/a\epsilon^{-(d+2)})$
Table 1: Complexity of sampling algorithms. We denote by $x^{*}$ a global minimizer of $V$, by $V^{*}=V(x^{*})$ the global minimum of $V$, by $m_{2}$ the second moment of the target, by $p_{t}$ the density of the forward process (see Sec. 4), and $\|\cdot\|$ the operator norm for matrices. In “RD + ULA,” RD refers to a Reverse Diffusion algorithm and ULA to how the intermediate scores are estimated. Even though originally stated in KL for our work and the one of He et al. [2024], all the complexities are w.r.t. the Total Variation distance (obtained via the Pinsker inequality).
Our algorithm addresses the three limitations outlined in the sections above. When the dimension $d$ is fixed, we obtain a polynomial query complexity. This guarantee does not assume log-smoothness and applies to general Gaussian mixtures, as will be shown in the next subsection. Moreover, this guarantee does not require running the algorithm with any explicit knowledge of the target distribution’s constants $a,b,\beta$.
Overall comparison
We summarize in Table 1 how our algorithm compares to previous approaches in terms of assumptions, oracles required to run the algorithm and resulting complexity. Along with the work of He and Zhang [2025], our algorithm is the only one that is polynomial in the parameters of the distribution when $d$ is fixed. Furthermore, while our dissipativity assumption is stronger than finite second moment, we relax the log-smoothness assumption by semi-log-convexity which notably covers general Gaussian mixtures. Finally, as mentioned above, because their algorithm requires an $\epsilon$-approximation of the global minimum of $V$, they require an explicit upper-bound on the second order moment of $\mu$ which may not be available in practice.
Numerical illustration
In Figure 2, we consider a standard task in the literature: sampling from a mixture of 16 equally weighted Gaussians with unit variance and centers uniformly distributed in $40,40]^{2} [$Midgley et al., 2023]. We compare our algorithm against Unadjusted Langevin Algorithm (ULA) and to the reverse diffusion algorithm of Huang et al. [2024a] (RDMC). We also implemented the zeroth-order method of He et al. [2024], but it failed to converge. ULA was initialized from $\mathcal{N}(0,I_{2})$ and run for $5\times 10^{4}$ steps. All three methods used the same discretization step size $h=0.01$. Our reverse diffusion algorithm and RDMC were both run with $500$ reverse diffusion steps and both used $100$ samples to estimate the intermediate scores. As discussed in Sec. 4, while our samples are simply drawn from a Gaussian distribution, the samples used in RDMC are drawn from a auxiliary, multi-modal distribution generated via an inner ULA step. In this experiment, this inner ULA was initialzed with a standard Gaussian and was run for $100$ steps; in particular, we emphasize that while ULA and our algorithm were roughly given the same computational budget, the one given to RDMC was a hundred times superior. As a result, ULA and our algorithm took approximately one minute to run on a computer locally and the RDMC method required over an hour. We observe that unlike the two others, our algorithm successfully recovers all the modes.
Figure 1: Error in Wasserstein distance as a function of the between-mode distance.
In Figure 1, we monitor the convergence of the same three algorithms as a function of the problem difficulty, measured by the distance between the modes. Here, the target distribution is a mixture of three Gaussians in two dimensions. It has equal weights $1/3$, equidistant modes located at a distance $R$ from the origin, and different covariances $(I,I/2,I/4)$. The final error is measured in Wasserstein distance, because it can be easily approximated using samples: we use $500$ samples generated by the sampling algorithm and $500$ samples from the true target distribution. We allowed each of the three algorithms $10^{6}$ queries to the potential $V$ or to its gradient: we performed $10^{6}$ iterations for ULA, we used $100$ steps of reverse diffusion for our algorithm and RDMC. We used $100$ inner steps of ULA for RDMC that was initialized with a standard Gaussian. We used $100$ particles for score estimation for RDMC and $10000$ for our algorithm. As expected, as the between-mode distance grows with $R$, our algorithm yields the lowest error.
Figure 2: From left to right: our algorithm vs. ULA vs. Huang et al. [2024a]. The color scheme indicates the probability density value of the distribution we want to sample from (dark is low probability density, bright is high probability). The blue dots are the samples produced by the algorithm.
3.2 Application to Gaussian mixtures
We now apply our results to derive provable sampling guarantees for general Gaussian mixtures. Apart from the very recent exception of Lytras and Mertikopoulos [2025] that we discuss below, note that despite their wide popularity, no sampling guarantees were yet derived for general Gaussian mixtures. Indeed, unless they verify some specific assumptions such as identical covariance matrices among components [Cordero-Encinar et al., 2025, Lemma B.1], general Gaussian mixtures do not satisfy the log-smoothness assumption, thus they do not fit the framework of many previously discussed works. In fact, their gradient may not even be Hölder continuous: consider the simple counter-example of a two-dimensional mixture $\mu=0.5\mathcal{N}(0,\Sigma_{1})+0.5\mathcal{N}(0,\Sigma_{2})$ with covariances $\Sigma_{1}=diag(1,0.5)$ and $\Sigma_{2}=diag(0.5,1)$. On the diagonal $x=y$, the score is $\nabla\log(\mu)(x,x)=-3/2(x,x)$, while near the diagonal, right above it for instance, the score behaves asymptotically as $\nabla\log(\mu)(x,x)\sim_{x\to+\infty}-(2x,x)$; we provide a rigorous analysis in Appendix B.2. Fortunately though, Gaussian mixtures do verify Assumptions 1-2, as we next show.
Proposition 2
Let $\mu=\sum_{i=1}^{p}w_{i}\mathcal{N}(\mu_{i},\Sigma_{i})$ and denote $\lambda_{min}>0$ (resp. $\lambda_{max}$) the minimum (resp. maximum) eigenvalue of the covariance matrices $\Sigma_{i}$. It holds that $\nabla^{2}\log(\mu)\preceq I_{d}/\lambda_{min}$ and that for all $x\in\mathbb{R}^{d}$, $\langle-\nabla\log(\mu)(x),x\rangle\geq\|x\|^{2}/(2\lambda_{max})-\lambda_{max}\max_{i}(\|\mu_{i}\|/\lambda_{min})^{2}$.
The proof is deferred to Appendix B.3. Combined with Theorem 1, Proposition 2 shows that we can sample any Gaussian mixture with average precision $\epsilon$ in $\mathop{\mathrm{KL}}\nolimits$ in $O(\mathrm{poly}(\kappa,R,d,\lambda_{min}^{-d},\epsilon^{-d}))$ queries to $V$ where $R=\max_{i}\|\mu_{i}\|$ and $\kappa=\lambda_{max}/\lambda_{min}$. There exists a relatively recent literature seeking to relax the log-smoothness assumption. For instance,  Chatterji et al. [2019], Erdogdu and Hosseinzadeh [2021], Nguyen et al. [2021] work under weak smoothness, i.e. $\alpha$-Hölder continuous gradient of the potential for $\alpha$ in $[0,1]$ (recall that $\alpha$=1 recovers the smooth case). However, as shown above, this relaxation is not sufficient yet to cover general Gaussian mixtures. The only reference we know of that does is the work of Lytras and Mertikopoulos [2025], who used a regularized version of Langevin to relax the global smoothness condition with local Lipschitz smoothness and polynomial growth of the Lipschitz constant. Yet crucially, their complexity bound scales as a polynomial of the Poincaré constant of $\mu$. We prove in Appendix B.4 that for the mixture $\mu=0.5\mathcal{N}(Ru,\lambda_{\max}I_{d})+0.5\mathcal{N}(-Ru,\lambda_{\min}I_{d})$ with $u$ any unit vector and $0<\lambda_{\min}\leq\lambda_{\max}$, this constant is at least $(R^{2}e^{R^{2}/(2\lambda_{\max})})/2$. In particular, this method still degrades exponentially with the multi-modality parameter $R$.
4 Our sampling algorithm
In this section, we introduce the reverse diffusion framework and explain how it reduces the sampling problem to that of estimating the scores along the forward Ornstein-Uhlenbeck (OU) process.
@@ 4.1 Reverse diffusion: from sampling to score estimation
Reverse diffusion methods emerged as an alternative to Langevin-based samplers in order to overcome metastability and were first introduced to the ML community in Song et al. [2021]. They rely on the so-called forward process
[…]
which corresponds to the standard OU process initialized at $\mu$, that is a specific case of a Langevin diffusion targeting a standard Gaussian, that we will denote $\pi$. Note that since the target of this process, the standard Gaussian, is $1$-strongly log-concave, the resulting process converges exponentially fast to the equilibrium. In order to sample from $\mu$, reverse diffusion algorithms rely on the semi-discretized backward process: given a horizon $T$ that we discretize as $0=t_{0}\leq t_{1}\leq\cdots t_{N-1}\leq t_{N}=T$, the latter writes
@@
with $Y_{0}\sim\pi$ and where all iterations can be solved in closed form. Because the forward process Eq. 1 converges exponentially fast, we can expect the initialization error $Y_{0}\sim\pi$ instead of $Y_{0}\sim p_{T}$ to be small after a short time $T$. Furthermore, if the proxies $s_{t_{k}}$ are sufficiently accurate, one can expect that the process output by the approximate scheme Eq. 3 has a distribution that is close to the target $\mu$. Over the past three years, several works provided quantitative bounds of the error induced by the discretization, the use of an approximate score and the initialization error with respect to different divergences and under various assumptions [Bortoli et al., 2021, Lee et al., 2022, Chen et al., 2023b, a, Conforti et al., 2025]. Yet, we shall rely exclusively on the following theorem as it is the most suited to our framework.
Theorem 3 (Conforti et al. [2025])
Assume that $\mu\propto e^{-V}$ has finite Fisher-information w.r.t. $\pi$ the standard gaussian density in $\mathbb{R}^{d}$:
$\mathcal{I}(\mu,\pi)\coloneqq\int\|x-\nabla V(x)\|^{2}\mathrm{d}\mu(x)<+\infty\,.$
Then, for the constant step-size discretization $t_{k}=kT/N$, denoting $p$ the distribution of the sample $Y_{T}$ output by Eq. 3, it holds that
$\mathop{\mathrm{KL}}\nolimits(\mu,p)\lesssim(d+m_{2})e^{-T}+\frac{1}{N}\sum_{k=1}^{N}\|\nabla\log(p_{t_{k}})-s_{t_{k}}\|_{L^{2}(p_{t_{k}})}^{2}+\frac{T}{N}\mathcal{I}(\mu,\pi)\,$,
@@
where $m_{2}$ is the second order moment of $\mu$ and where $\lesssim$ hides a universal constant.
@@
The previous theorem shows that under mild assumptions that notably allow for multi-modality, the problem of sampling from $\mu$ can be transferred into a score approximation problem along the forward process. In the next subsection, we present an estimator for these intermediate scores that is tractable given the knowledge of the unnormalized density $\mu\propto e^{-V}$.
4.2 Derivation of an estimator
The key observation to derive an estimator of the intermediate scores is that the forward process Eq. 1 is nearly available in closed form. However, this closed form may be written in different manners.
Different expressions of the scores
Consider Eq. 1 integrates to $X_{t}=\sqrt{\lambda_{t}}X_{0}+\sqrt{1-\lambda_{t}}Z\,$, where $\lambda_{t}=e^{-2t}$, $Z$ is a standard $d$-dimensional Gaussian, and $X_{0}$ is a random variable simulating the target distribution. The corresponding density convolves the target density with a Gaussian, as
$p_{t}^{V}(x)$ $=\frac{1}{\sqrt{1-\lambda_{t}}}\pi\left(\frac{x}{\sqrt{1-\lambda_{t}}}\right)\ast\frac{1}{\sqrt{\lambda_{t}}}\mu\left(\frac{x}{\sqrt{\lambda_{t}}}\right)\,$, (4)
where $\pi$ is a standard Gaussian and $\mu(\cdot)\propto\exp(-V(\cdot))$ is the target distribution. From this expression, we can obtain different formulas for the scores. We retain
  1. 1.
    $\nabla\log p_{t}^{V}(x)=\frac{1}{\sqrt{\lambda_{t}}}\mathbb{E}_{y|x}\left[\nabla\log\mu\left(\frac{x-y}{\sqrt{\lambda_{t}}}\right)\right],\quad y|x\sim\propto\mu\left(\frac{x-y}{\sqrt{\lambda_{t}}}\right)\times\mathcal{N}(y;0,(1-\lambda_{t})I)$   ,
  2. 2.
    $\nabla\log p_{t}^{V}(x)=\frac{1}{\sqrt{\lambda_{t}}}\frac{\mathbb{E}_{y}\left[\nabla\log\mu\left(\frac{x-\sqrt{1-\lambda_{t}}y}{\sqrt{\lambda_{t}}}\right)\mu\left(\frac{x-\sqrt{1-\lambda_{t}}y}{\sqrt{\lambda_{t}}}\right)\right]}{\mathbb{E}_{y}[\mu\left(\frac{x-\sqrt{1-\lambda_{t}}y}{\sqrt{\lambda_{t}}}\right)]},\quad y\sim\mathcal{N}(y;0,I)$   ,
  3. 3.
    $\nabla\log p_{t}^{V}(x)=\frac{1}{1-\lambda_{t}}\left(\mathbb{E}_{y|x}[y]-x\right),\quad y|x\sim\mu\left(\frac{y}{\sqrt{\lambda_{t}}}\right)\times\mathcal{N}(y;x,(1-\lambda_{t})I)$   ,
  4. 4.
    $\nabla\log p_{t}^{V}(x)=\frac{1}{\sqrt{1-\lambda_{t}}}\frac{\mathbb{E}_{y}\left[y\mu\left(\frac{x+\sqrt{1-\lambda_{t}}y}{\sqrt{\lambda_{t}}}\right)\right]}{\mathbb{E}_{y}\left[\mu\left(\frac{x+\sqrt{1-\lambda_{t}}y}{\sqrt{\lambda_{t}}}\right)\right]},\quad y\sim\mathcal{N}(y;0,I)\,.$
Derivations can be found in Appendix A. The second [Akhound-Sadegh et al., 2024], third [Huang et al., 2024a, He et al., 2024, Grenioux et al., 2024], and fourth [Saremi et al., 2024] identities have been used to build Monte Carlo estimators of the score. Hence, one natural way to group these identities is by how difficult it is to draw the samples $y$. The second and fourth identities sample from a Gaussian distribution, whereas the first and third identities sample from an auxiliary multi-modal distribution from generating $y|x$. As $t$ gets closer to zero, this auxiliary distribution resembles the original target, so sampling from it progressively becomes as difficult as sampling from the original target density. This observation explains why, unsurprisingly, Huang et al. [2024a], He et al. [2024] do not improve upon existing results. Another way to group these identities is whether or not they require evaluating the score of the target distribution: identities one and two do and are referred to as TSI estimators, referring to the Target Score Identity (TSI) used to obtain them [Bortoli et al., 2024]; identities three and four do not and are referred to as DSI estimators, referring to a Denoising Score Identity (DSI) used to obtain them [He et al., 2025]. Combinations of TSI and DSI estimators have also been considered [Phillips et al., 2024, He et al., 2025].
A self-normalized estimator
We now focus on the estimator of the scores that we use. It is a rewrite of the fourth identity and is a ratio of expectations under Gaussians
$\nabla\log(p_{t}^{V})(z)=\frac{-1}{1-e^{-2t}}\frac{\mathbb{E}[Y_{t}e^{-V(e^{t}(z-Y_{t}))}]}{\mathbb{E}[e^{-V(e^{t}(z-Y_{t}))}]}\,$,
with $Y_{t}\sim\mathcal{N}(0,(1-e^{-2t})I_{d})$ which is easy to simulate. While conventional statistical wisdom may suggest using independent samples to estimate both the numerator and the denominator, we voluntarily choose to correlate them and implement instead